OpenStudy (anonymous):

center of circle is (-2,1) and one point is (-5,3) what is the diameter

OpenStudy (cuanchi):

\[5-2=3; 3-1=2 ; r^{2}=3^{2}+2^{2}; \rightarrow r=\sqrt{13}\]